# On the Adversarial Robustness of Learning-based Conformal Novelty Detection

### v3

@@ I Introduction
How to incorporate existing adversarial machine learning algorithms? We propose a general methodology to make existing attack algorithms more effective. As a proof of concept, we adopt two popular off-the-shelf black-box adversarial attack algorithms, HopSkipJump [15] and Boundary Attack [16], which only require the predicted label for refined data perturbation (see a detailed discussion in Section III-A). These algorithms allow us to effectively change the score for the decision-making while minimizing the required changes in data. This approach enables directly attacks on the raw data, which is much more realistic than attacking scores of the data after some fixed transformation (e.g., p-values). For instance, there are works that focus on attacking p-values directly in the widely used Benjamini-Hochberg (BH) procedure for FDR control: a distributed setting using the BH procedure in the presence of compromised Byzantines [17], and on the adversarial robustness of the BH procedure through perturbation of p-values [18]. For the latter work by Chen et al. [18], see a detailed comparison between their INCREASE-c method and our approach in Section E to demonstrate the perturbations required for effective attacks.
@@ I-A Contributions and Outline
The main contributions of this work are threefold. First, we introduce and design the adversarial attacks of the AdaDetect framework by proposing an oracle setting, when the attacker has access to both the model and test data labels, to quantify the upper bound on the loss in FDR (Section III-A). Second, our oracle setting naturally motivates the design of a practical query-based attack scheme (detailed in Section III-C), called the surrogate decision-based attack, where the attacker can only query for the labels of the test data. This works for both AdaDetect and the one by Bates et al. [4]. Third, in Section IV, the vulnerability of AdaDetect and Bates et al. [4] under our proposed attack strategies is extensively evaluated using two popular and complementary adversarial machine learning attack algorithms: HopSkipJump and Boundary Attack. The code can be found in the supplementary material.
@@ II Background
where $R\vee 1:=\max\{R,1\}$. The power measures the detection performance of non-nulls, defined as $\text{power}=\mathbb{E}[(R-V)/(m_{1}\vee 1)]$, where $m_{1}=|\mathcal{H}_{1}|$ is the number of non-nulls in the test set.
@@ II-A The AdaDetect Scheme [7]
$\begin{split}&(Z_{1},\ldots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0})\mid(Z_{n+j}:j\in\mathcal{H}_{1})\overset{d}{=}\\ &(Z_{\pi(1)},\ldots,Z_{\pi(n)},Z_{\pi(n+i)}:i\in\mathcal{H}_{0})\mid(Z_{n+j}:j\in\mathcal{H}_{1})\end{split}$
@@
Step 1: Learn score function. Partition the data into the training data $\{Z_{i}\}_{i=1}^{k}$ and the mixed sample $\{Z_{i}\}_{i=k+1}^{n+m}$, which contains both calibration samples from $P_{0}$ and unlabeled test samples. Under the positive-unlabeled (PU) learning framework, apply a machine learning algorithm to learn a data-driven and measurable score function: $s:\mathcal{Z}\times\mathcal{Z}^{k}\times\mathcal{Z}^{n+m-k}\to\mathbb{R}$ as
@@
satisfying the following permutation invariance property. For any permutation $\pi$ of $\{k+1,\ldots,n+m\}$,
$s(z;(z_{1},\ldots,z_{k}),(z_{\pi(k+1)},\ldots,z_{\pi(n+m)}))=$
$\quad\quad\quad\quad s(z;(z_{1},\ldots,z_{k}),(z_{k+1},\ldots,z_{n+m})).$ (3)
@@
$\begin{split}&O_{i}=s(Z_{i};(Z_{1},\ldots,Z_{k}),(Z_{k+1},\ldots,Z_{n+m})),\\ &\quad i\in[k+1:n+m].\end{split}$ (4)
@@
In Section IV, we carry out experiments under both the marginal p-value setting and the CCV setting.
@@ III-A Oracle Setting: Worst-case Attack Scheme
where $Y^{*}_{i}=0$ for $i\in\mathcal{H}_{0}$, $Y^{*}_{i}=1$ for $i\in\mathcal{H}_{1}$ and then train a score function
@@
$:=f_{\text{attack}}(Z_{n+i};\,\{Z_{1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\},$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad(Z_{n+j}:j\in A\cup\mathcal{H}_{1}))$ (7)
such that $\mathbf{1}\{g(Z_{n+i})\geq 0.5\}\neq\mathbf{1}\{g(\widetilde{Z}_{n+i})\geq 0.5\}$, meaning that the decision is altered. Here we write
@@
as an unordered set to highlight that $f_{\text{attack}}$ does not depend on the order of elements in this set.
Step 3: Applying AdaDetect on the attacked data. After the attack, the user applies AdaDetect and computes the score function as the first step. As the data is now changed by the attacker, we denote the score function after the attack by $\tilde{s}$, and the empirical $p$-values after the attack by ${\tilde{p}}_{i}$ for $i\in[1:m]$. We stress that $\tilde{s}$ still satisfies equation II-A.
@@
$f_{\text{attack}}$ does not depend on the order of elements in $\{Z_{1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$.
In Section IV, we evaluate adversarial robustness using two representative decision-based attacks: the Boundary Attack [16] and the HopSkipJumpAttack (HSJA) [15]. Decision-based attacks, which rely solely on query access to the decision function, capture adversarial capabilities more realistically than white-box or score-based methods. Moreover, evaluating FDR loss under such attacks provides a stringent robustness assessment, as these “blind” perturbations often induce more severe failures than those observed under other threat models. The two algorithms were chosen for their complementary properties. The Boundary Attack is a seminal decision-based approach that operates via a random-walk strategy, starting from an adversarial example and progressively reducing the perturbation while remaining misclassified. It is conceptually simple, model-agnostic, and widely adopted as a baseline in the literature. In contrast, HSJA is a more recent attack that achieves state-of-the-art query efficiency by combining binary search with adaptive estimation of the decision boundary’s normal vector. While both attacks require only hard-label access to the model, Boundary Attack provides a robust baseline, whereas HSJA represents a stronger and more query-efficient adversary. Together, they allow us to assess robustness against both classical and modern decision-based adversarial paradigms.
The attack in the last step pushes the score above the decision boundary. It is important to note that the output of the HSJA scheme [15], i.e., $f_{\text{attack}}=\text{HSJA}$ indeed does not depend on the order of elements in $\{Z_{1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}\}$. The same holds for the Boundary Attack [16].
@@ III-B Analysis
In our oracle setting, we denote the corresponding FDR as $\text{FDR}^{*}_{\text{attack}}$. Our main theorem quantifies the loss in FDR caused by the attack. The proof is deferred to Appendix C.
@@
Consider that $\mathcal{A}$ is a fixed set of indices with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation II-A and the attack scheme $f_{attack}$ being order-invariant as in equation 7, the FDR after the attack is
$\text{FDR}^{*}_{\text{attack}}\leq\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}_{m_{a}}\lor 1}\right],$ (8)
@@
It turns out that the proof techniques for this oracle setting in Theorem 1 can be adapted to less stringent settings where the true labels of test samples are unknown to the attacker. We report them in Appendix D. In the next section, we propose a heuristic algorithm that is motivated by our oracle setting.
The key lemma below shows the conditional exchangeability we need for Theorem 1. First, we introduce the following notation for simplicity of presentation. Denote the number of unattacked true null test samples by $\tilde{m}_{0}$. In order to simplify notation, let
$U_{\setminus\mathcal{A}}$ $=(U_{1},\dots,U_{n+\tilde{m}_{0}}):=$
$\quad\quad\quad\quad(Z_{1},\dots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}),$
@@
With a slight abuse of notation, the condition equation 7 on $f_{\text{attack}}$ can be simplified as
@@
where we stress that $f_{\text{attack}}$ does not depend on the order of the elements in $U_{\setminus\mathcal{A}}$. The following lemma establishes a crucial property not addressed in the original AdaDetect:
@@
The key insight is that by conditioning on not only non-null data $V$ but also the attack outcomes $\widetilde{U}_{\mathcal{A}}$, the unattacked null samples maintain their exchangeable structure. See the proof in Appendix A.
@@
$\begin{split}&(U_{k+1},\dots,U_{n+\tilde{m}_{0}})\mid V\cup\widetilde{U}_{\mathcal{A}}\overset{d}{=}\\ &\quad\quad\quad\quad\quad(U_{\pi(k+1)},\dots,U_{\pi(n+\tilde{m}_{0})})\mid V\cup\widetilde{U}_{\mathcal{A}}\end{split}$
@@
Remark 4.
For our theoretical analysis, we find it is sufficient to establish exchangeability for the unattacked true null elements indexed from $k+1$ onwards to derive the FDR upper bound. While the original AdaDetect analysis demonstrates exchangeability for all true null elements including the first $k$ training samples, this broader exchangeability does not hold in our adversarial setting due to the attack’s dependence on the complete dataset. Specifically, their exchangeability expression covers $(Z_{1},\ldots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0})$ conditioned on $V$, while our restricted result only requires $(Z_{k+1},\ldots,$ $Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A})$ to be exchangeable conditioned on $\widetilde{U}_{\mathcal{A}}$ and $V$.
Now we show that the conditional exchangeability of data in Lemma 1 can be carried over to the scores. See the proof in Appendix B.
@@
$(\tilde{s}(U_{k+1}),\ldots,\tilde{s}(U_{n+\tilde{m}_{0}}))\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup$
$(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})\overset{d}{=}(\tilde{s}(U_{\pi(k+1)}),\ldots,\tilde{s}(U_{\pi(n+\tilde{m}_{0})}))\mid$
$\quad\quad\quad\quad\quad(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$
@@
Building on Lemma 1 and Lemma 2, the following result follows directly from Theorem A.1 (iii) and (iv) from [7] and we skip the proof.
@@
$\widetilde{W}_{i}\;=\;$ $(\{S_{k+1},\dots,S_{n},S_{n+i}\}\;,\;$
$\quad(S_{n+j}:j\neq i,j\in\mathcal{H}_{0})\;,(S_{n+j}:j\in\mathcal{H}_{1})\,).$

### v4

@@ I Introduction
How to incorporate existing adversarial machine learning algorithms? We propose a general methodology to make existing attack algorithms more effective. As a proof of concept, we adopt two popular off-the-shelf black-box adversarial attack algorithms, HopSkipJump [15] and Boundary Attack [16], which only require the predicted label for refined data perturbation (see a detailed discussion in Section III-A). These algorithms allow us to effectively change the score for the decision-making while minimizing the required changes in data. This approach enables directly attacks on the raw data, which is much more realistic than attacking scores of the data after some fixed transformation (e.g., $p$-values). For instance, there are works that focus on attacking $p$-values directly in the widely used Benjamini-Hochberg (BH) procedure for FDR control: a distributed setting using the BH procedure in the presence of compromised Byzantines [17], and on the adversarial robustness of the BH procedure through perturbation of $p$-values [18]. For the seminal work by Chen et al. [18], see Appendix C for a detailed comparison between their INCREASE-c method and our approach in to demonstrate the perturbations required for effective attacks.
@@ I-A Contributions and Outline
The main contributions of this work are threefold. First, we formulate and design adversarial attacks of the AdaDetect framework by proposing an oracle setting, when the attacker has access to both the model and test data labels, to quantify the upper bound on the loss in FDR (Section III-A). Second, our oracle setting naturally motivates the design of a practical query-based attack scheme (Section III-C), called the surrogate decision-based attack, where the attacker can only query for the labels of the test data. This works for both AdaDetect and the one by Bates et al. [4]. Third, in Section IV, the vulnerability of AdaDetect and Bates et al. [4] under our proposed attack strategies is extensively evaluated using two popular and complementary adversarial machine learning attack algorithms: HopSkipJump and Boundary Attack. The code can be found in the supplementary material.
@@ II Background
where $R\vee 1:=\max\{R,1\}$. The power measures the detection performance of non-nulls, defined as $\text{power}=\mathbb{E}[(R-V)/(m_{1}\vee 1)]$, where $m_{1}=|\mathcal{H}_{1}|$ is the number of non-nulls in the test set. We write $n_{1}:n_{2}]$, for $n_{1}<n_{2}$, to denote the set of consecutive numbers from $n_{1}$ to $n_{2}$, i.e., $\{n_{1},n_{1}+1,...,n_{2}\}$.
@@ II-A The AdaDetect Scheme [7]
$\begin{split}&(Z_{1},\ldots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0})\mid(Z_{n+j}:j\in\mathcal{H}_{1})\overset{d}{=}\\ &\quad\quad\quad\quad(Z_{\pi(1)},\ldots,Z_{\pi(n)},Z_{\pi(n+i)}:i\in\mathcal{H}_{0})\mid(Z_{n+j}:j\in\mathcal{H}_{1})\end{split}$
@@
Step 1: Learn score function. Partition the data into the training data $\{Z_{i}\}_{i=1}^{k}$ and the mixed sample $\{Z_{i}\}_{i=k+1}^{n+m}$, which contains both calibration samples from $P_{0}$ and unlabeled test samples. Under the positive-unlabeled (PU) learning framework (labeling the first $k$ samples as null and the rest of them as non-nulls), apply a machine learning algorithm to learn a data-driven and measurable score function: $s:\mathcal{Z}\times\mathcal{Z}^{k}\times\mathcal{Z}^{n+m-k}\to\mathbb{R}$ as
@@
satisfying the following permutation property. For any permutation $\pi$ of $\{k+1,\ldots,n+m\}$,
$s(z;(z_{1},\ldots,z_{k}),(z_{\pi(k+1)},\ldots,z_{\pi(n+m)}))=s(z;(z_{1},\ldots,z_{k}),(z_{k+1},\ldots,z_{n+m})).$ (3)
@@
$\begin{split}&O_{i}=s(Z_{i};(Z_{1},\ldots,Z_{k}),(Z_{k+1},\ldots,Z_{n+m})),i\in[k+1:n+m].\end{split}$ (4)
@@
In Section IV, we carry out experiments under both the marginal $p$-value and the CCV settings.
@@ III-A Oracle Setting: Worst-case Attack Scheme
where $Y^{*}_{i}=0$ for $i\in\mathcal{H}_{0}$, $Y^{*}_{i}=1$ for $i\in\mathcal{H}_{1}$ are the true labels. Then train a score function
@@
$:=f_{\text{attack}}(Z_{n+i}\,;\,\{Z_{1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\},(Z_{n+j}:j\in A\cup\mathcal{H}_{1}))$ (7)
such that $\mathbf{1}\{g(Z_{n+i})\geq 0.5\}\neq\mathbf{1}\{g(\widetilde{Z}_{n+i})\geq 0.5\}$, meaning that the decision is altered. We write
@@
as an unordered set to highlight that $f_{\text{attack}}$ does not depend on the order of elements in this set. In our experiments (see Section IV), we evaluate adversarial robustness using two representative decision-based attacks: the HopSkipJumpAttack (HSJA) [15] and the Boundary Attack [16]. Therefore, our attack scheme inherits their underlying optimized perturbation. In both HSJA and Boundary, they minimize $\lVert z-{\tilde{z}}\rVert^{2}$ such that the label is flipped, namely, $\mathbf{1}\{g(z)\geq 0.5\}\neq\mathbf{1}\{g(\tilde{z})\geq 0.5\}$.
Step 3: Apply AdaDetect on the contaminated (or attacked) data. After the attack, the user applies AdaDetect and computes the score function as the first step. As the data is now changed by the attacker, we denote the score function after the attack by $\tilde{s}(z)$, and the empirical $p$-values after the attack by ${\tilde{p}}_{i}$ for $i\in[1:m]$. We stress that $\tilde{s}(z)$ still satisfies equation 3. The number of rejections is denoted by $\widetilde{R}(m_{a})$, where we explicitly shows its dependence on $m_{a}$ for our main results in the next section.
@@
$f_{\text{attack}}(\cdot\,;\,g(z))$ does not depend on the order of elements in $\{Z_{1},...,Z_{n},Z_{n+j}:j\in\mathcal{H}_{0}\setminus\mathcal{A}\}$.
Decision-based attacks, including HSJA and Boundary attack, which rely solely on query access to the decision function, capture adversarial capabilities more realistically than white-box or score-based methods. Moreover, evaluating FDR loss under such attacks provides a stringent robustness assessment, as these “blind” perturbations often induce more severe failures than those observed under other threat models. The two algorithms were chosen for their complementary properties. The Boundary Attack is a seminal decision-based approach that operates via a random-walk strategy, starting from an adversarial example and progressively reducing the perturbation while remaining misclassified. It is conceptually simple, model-agnostic, and widely adopted as a baseline in the literature. In contrast, HSJA is a more recent attack that achieves state-of-the-art query efficiency by combining binary search with adaptive estimation of the decision boundary’s normal vector. While both attacks require only hard-label access to the model, Boundary Attack provides a robust baseline, whereas HSJA represents a stronger and more query-efficient adversary. Together, they allow us to assess robustness against both classical and modern decision-based adversarial paradigms.
The attack in the last step pushes the score above the decision boundary. It is important to note that the output of the HSJA scheme [15] indeed does not depend on the order of elements in $\{Z_{1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}\}$. The same holds for the Boundary Attack [16].
@@ III-B Analysis
In our oracle setting, we denote the corresponding FDR as $\text{FDR}^{*}_{\text{attack}}$. Our main theorem quantifies the loss in FDR caused by the attack. Recall that $\widetilde{R}(m_{a})$ denotes the number of rejections when the user applies Adadetect on the contaminated data samples. The proof is deferred to Appendix A.
@@
Consider that $\mathcal{A}$ is a fixed set of indices with $m_{a}=|\mathcal{A}|$. Under Assumption 1, with the score function $\tilde{s}$ satisfying the permutation invariance property in equation 3 and the attack scheme $f_{attack}$ being order-invariant as in equation 7, the FDR after the attack is
$\text{FDR}^{*}_{\text{attack}}\leq\frac{m_{0}-m_{a}}{m}\alpha+m_{a}\cdot\mathbb{E}\left[\frac{1}{{\widetilde{R}}(m_{a})\lor 1}\right],$ (8)
where the expectations are taken over the randomness in the training and test samples $\{Z_{j}\}_{j=1}^{m+n}$. Furthermore, we have $\text{FDR}^{*}_{\text{attack}}\leq\alpha$, as long as
$h(m_{a})\leq\frac{m_{1}}{m}\,\alpha\,,\quad\text{ where }\quad h(x):=x\,\cdot\biggl(\mathbb{E}\left[\frac{1}{{\widetilde{R}}(x)\lor 1}\right]-\frac{\alpha}{m}\biggr).$ (9)
It is straightforward to see that equation 9 comes from upper-bounding equation 8 by $\alpha$. As the bound on $m_{a}$ is implicit, we provide some clarifications to help gain a better understanding of the valid $m_{a}$’s that satisfies equation 9. Consider a simple Gaussian setting (see details in Section IV), for $m=1000$ with $m_{1}=100,200,300$, the corresponding ranges of $m_{a}$ are: $m_{a}=1$, $m_{a}\in[1:4]$, and $m_{a}\in[1:9]$, respectively. As a rough approximation, one can even upper bound ${\widetilde{R}}(m_{a})$ by $m$, which gives $m_{a}\leq\frac{\alpha}{1-\alpha}m_{1}$ (for instance, it gives $m_{a}\leq 11$ for $m_{1}=100$ and $\alpha=0.1$), but it is important to keep in mind that this upper bound on $m_{a}$ no longer guarantees $\text{FDR}^{*}_{\text{attack}}\leq\alpha$ and thus serves as a loose bound on the valid $m_{a}$ satisfying equation 9. These all indicate that when $m_{a}$ is very small, AdaDetect is robust to any adversarial attacks. On the other hand, as detailed in our numerical experiments (Section IV), we observe that $\text{FDR}^{*}_{\text{attack}}$ goes beyond $\alpha$ fairly quickly as $m_{a}$ grows.
@@
It turns out that the proof techniques for this oracle setting in Theorem 1 can be adapted to less stringent settings where the true labels of test samples are unknown to the attacker (see Appendix B).
The key lemmas below show the conditional exchangeability we need for Theorem 1. First, we introduce the following notation for simplicity of presentation. Denote the number of unattacked true null test samples by $\tilde{m}_{0}$. In order to simplify notation, let
$U_{\setminus\mathcal{A}}$ $=(U_{1},\dots,U_{n+\tilde{m}_{0}}):=(Z_{1},\dots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}),$
@@
With a slight abuse of notation, the condition equation 7 on $f_{\text{attack}}(\cdot\,;\,g(z))$ can be simplified as
@@
where we stress that $f_{\text{attack}}$ does not depend on the order of the elements in $U_{\setminus\mathcal{A}}$. Note that we have $Z_{n+i}=\widetilde{Z}_{n+i}$ for any $i\notin\mathcal{A}$. We will present two main lemmas that establish a crucial property that is unique to our adversarial setting:
@@
The key insight is that by conditioning on not only non-null data $V$ but also the attack outcomes $\widetilde{U}_{\mathcal{A}}$, the unattacked null samples maintain their exchangeable structure.
The main technical challenge arises from a certain form of “matching” between the invariance property (equation 7) of $f_{\text{attack}}(\cdot\,;\,g(z))$ by the attacker and the invariance property (equation 3) of score function $\tilde{s}(z)$ by the user on the contaminated data. (Recall from Step 3 of the oracle setting, $\tilde{s}$ satisfies equation 3 because of the PU classifier.) We explain this important point here before presenting the two lemmas. It turns out that for our analysis to work, we need to work with the exchangeability for the unattacked true null elements indexed from $k+1$ and onwards (i.e., $\{U_{k+1},\dots,U_{n+\tilde{m}_{0}}\}=\{Z_{k+1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}\}$) to derive an upper bound on FDR after attack. While the original AdaDetect analysis demonstrates exchangeability for all true null elements including the first $k$ training samples (i.e., $(Z_{1},\ldots,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0})$), this broader exchangeability does not hold in our adversarial setting due to the following subtle yet important consideration.
Both of the attack function $f_{\text{attack}}$ and the score function $\tilde{s}$ should be invariant w.r.t. $\{U_{k+1},\dots,U_{n+\tilde{m}_{0}}\}=\{Z_{k+1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A}\}$.
Fortunately, this holds because $g(z)$ is trained on true labels (assigning the same label to this set of data samples) and $f_{\text{attack}}(\cdot\,;\,g(z))$ depends on the data through $g(z)$, while $\tilde{s}(z)$ uses PU which also assigns the same label to this set, since PU labels the first $k$ samples $\{Z_{i}\}_{i=1}^{k}$ by 0 and the rest of the samples $\{Z_{i}\}_{i=k+1}^{n+m}$ by $1$.
@@
$\begin{split}&(U_{k+1},\dots,U_{n+\tilde{m}_{0}})\mid V\cup\widetilde{U}_{\mathcal{A}}\overset{d}{=}(U_{\pi(k+1)},\dots,U_{\pi(n+\tilde{m}_{0})})\mid V\cup\widetilde{U}_{\mathcal{A}}\end{split}$
@@
Proof.
and we write this in a compact form as $\widetilde{U}_{\mathcal{A}}:=f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)$. We want to show that
According to Proposition 1, $f_{\text{attack}}$ does not depend on the order of the elements from $U_{k+1}$ to $U_{n+\tilde{m}_{0}}$ in $U_{\setminus\mathcal{A}}$, so we have $\widetilde{U}_{\mathcal{A}}=f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V)$. For any measurable set $\mathcal{U}_{\setminus\mathcal{A}}$ in the support of $U_{\setminus\mathcal{A}}$, we have
$\mathrm{P}(U_{\setminus\mathcal{A}}\in\mathcal{U}_{\setminus\mathcal{A}},f_{\text{attack}}(U_{\mathcal{A}};U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}|V=v)$
$\quad\quad\quad=\mathrm{P}(U_{\setminus\mathcal{A}}^{\pi}\in\mathcal{U}_{\setminus\mathcal{A}},f_{\text{attack}}(U_{\mathcal{A}};U_{\setminus\mathcal{A}}^{\pi},U_{\mathcal{A}}\cup V)=\tilde{u}_{\mathcal{A}}|V=v).$
Therefore,
Now we show that the conditional exchangeability of data in Lemma 1 can be carried over to the scores, building on the key property that both of $f_{\text{attack}}$ and $\tilde{s}$ are invariant w.r.t. $\{U_{k+1},\dots,U_{n+\tilde{m}_{0}}\}=\{Z_{k+1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus A\}$, as we alluded to before.
@@
$(\tilde{s}(U_{k+1}),\ldots,\tilde{s}(U_{n+\tilde{m}_{0}}))\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$
$\quad\quad\quad\quad\overset{d}{=}(\tilde{s}(U_{\pi(k+1)}),\ldots,\tilde{s}(U_{\pi(n+\tilde{m}_{0})}))\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$
@@
Proof.
First, we introduce the following notion $Q$ to capture not only the invariance property in equation 3 satisfied by $\tilde{s}$, but also the invariance property in equation 7 of $f_{\text{attack}}$,
$Q$ $:=h(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)$
$\overset{(a)}{=}((Z_{1},\dots,Z_{k}),\{Z_{k+1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus A\},(Z_{n+i},i\in\mathcal{H}_{1}\cup\mathcal{A})\,),$ (15)
where $(a)$ follows from (1) the key property that $f_{\text{attack}}$ and $\tilde{s}$ are invariant w.r.t. $\{Z_{k+1},...,Z_{n},Z_{n+i}:i\in\mathcal{H}_{0}\setminus A\}$, and (2) recall that $\widetilde{U}_{\mathcal{A}}:=f_{\text{attack}}(U_{\mathcal{A}}\,;\,U_{\setminus\mathcal{A}},U_{\mathcal{A}}\cup V)$. By Lemma 1, we have
Since $Q$ is a function of $(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)$, we have
Since $\pi$ keeps the first $k$ indices fixed, by the definition of $Q$, we have $h(U_{\setminus\mathcal{A}}^{\pi},\widetilde{U}_{\mathcal{A}},V)=h(U_{\setminus\mathcal{A}},\widetilde{U}_{\mathcal{A}},V)=Q$. Thus
Observe that the score terms $(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$ are deterministic functions of the conditional variables $\widetilde{U}_{\mathcal{A}},V,Q$; recall that $Z_{n+i}=\widetilde{Z}_{n+i}$ for any $i\notin\mathcal{A}$. Since $\pi(i)=i$ for all $i\leq k$, the exchangeability is preserved as follows,
$(S_{k+1},\dots,S_{n},S_{n+i}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A})$
$\quad\quad\overset{d}{=}(S_{\pi(k+1)},\dots,S_{\pi(n)},S_{\pi(n+i)}:i\in\mathcal{H}_{0}\setminus\mathcal{A})\mid(\tilde{s}(Z_{n+j}):j\in\mathcal{H}_{1})\cup(\tilde{s}(\widetilde{Z}_{n+j}):j\in\mathcal{A}).$
Building on Lemma 1 and Lemma 2, the following result follows directly from [7, Theorem A.1 (iii) and (iv)] and we skip the proof.
@@
$\widetilde{W}_{i}\;=\;$ $(\{S_{k+1},\dots,S_{n},S_{n+i}\},\;(S_{n+j}:j\neq i,j\in\mathcal{H}_{0}\setminus\mathcal{A})\;,\;$
$\hskip 120.00018pt(S_{n+j}:j\in\mathcal{H}_{1})\,\cup(S_{n+j}:j\neq i,j\in\mathcal{A}).$